# Enhancing Diversity in Parallel Agents: A Maximum State Entropy Exploration Story

### v1

@@ 2 Agents1 Agent $K^{\prime}=2K$6 Agents1 Agent $K^{\prime}=6K$
(a) State Entropy (ROOMDET)
(b) State Entropy (ROOMS)
(c) State Entropy (LABDET)
(d) State Entropy (LABSTOCH)
(e) Support (ROOMDET)
(f) Support (ROOMS)
(g) Support (LABDET)
(h) Support (LABSTOCH)
@@
  • Policy Gradient Optimization – This section analyzes the results of the learning process based on Algorithm 1, evaluating the performance of the learned policy in terms of normalized entropy and support size.
  • Dataset Entropy Evaluation – This section takes a deeper look at the entropy of a dataset collected from the policies trained as in the previous section.
  • Offline RL – This section provides a comparative analysis of how the dataset collected with parallel maximum entropy agents benefits the performance of offline RL.
For all the experiments, we consider variations of discrete gridworld domains, which we depicted in Appendix B (see Figures 66). One family (ROOM) is composed of two rooms connected by a narrow corridor, with a total of 43 states and 4 actions (one to move in each direction). The other (LAB) is a maze of various connected paths, with same number of total states and actions as before. For each family, we consider a deterministic version and a stochastic version in which the agent’s chosen actions are flipped to random actions with some probability. Appendix C reports experiments details with further visualizations and results. The code to reproduce the experiments is made available at the following link.
@@
We test the ability of our algorithm, PGPSE (Algorithm 1), in training parallel agents for maximum state entropy exploration. We run multiple instances of the algorithm with an increasing number of agents $m\in\{2,4,6\}$. [^3] In the main text we omit the result with 4 agents, which are reported in Figure 7 of the Appendix. As a baseline, we consider the performance of the same algorithm training a single agent for the finite-trial objective (Eq. 2), with a number of trials $K^{\prime}$ matching the number of trajectories taken by the parallel agents. We compare the results in terms of (normalized) state entropy and the size of the entropy support (i.e., the number of unique visited states) in the ROOM, LAB domains, deterministic and stochastic versions.
Figure 1 report the resulting learning curves, obtained over 5 independent runs. For a fair comparison of algorithms with varying number of learners, the performance is reported as a function of the policy updates, which translate to $|\mathcal{B}|\times m\times|K|\times T$ total interactions with the environment. Across the board, the performance of parallel agents (solid lines) consistently outperforms the single agent counterparts (dashed lines). They achieves higher entropy values, thanks to their ability to explore more states as it is evident from the larger size of the support. This behavior can be attributed to the intrinsic diversity introduced by parallel agents: To optimize the team state distribution, each agent learns an individual behavior that is synergic with the others.
@@
Here we consider the policies trained as in the previous section to collect a single dataset, to better capture the capabilities of the respective policies after training. Figure 2 illustrates the normalized empirical state entropy of each dataset collected by the $m$ agents, compared against the entropy of the dataset collected by a single agent over $K=m$ trajectories (trained to maximize the entropy on the corresponding number of trajectories), and a random policy.
To construct these datasets, we sample one trajectory per agent from the parallel policies, evaluating them across 5 seeds. This procedure reveals that datasets generated through parallel exploration exhibit higher entropy than those collected by a single agent or a random policy. It is noteworthy that experience collected in parallel also exhibits minimal variance across runs, indicating that agents tend to generate trajectories in a quasi-deterministic manner, preserving consistency in the collected data.
2 Agents1 Agent $K^{\prime}=2K$6 Agents1 Agent $K^{\prime}=6K$
(a) Dataset Entropy (ROOMDET)
(b) Dataset Entropy (ROOMS)
(c) Dataset Entropy (LABDET)
(d) Dataset Entropy (LABSTOCH)
Figure 2: Entropy evaluation of the dataset generated by sampling from parallel policies, compared against both random and single-baselines across different environments. With the notation $K^{\prime}=mK$ we denoted the set in which the single agent had access to a multiple $m$ of trajectories available to the parallel agents. We report average and standard deviation over 5 inependent runs.
@@
Can datasets collected with parallel maximum entropy agents benefit offline RL algorithms? One practical application of the policies learn with PGPSE, in line with the motivation of this work, is to utilize the experience gained during the exploration phase to construct informative datasets for offline RL (Yarats et al., 2022).
Starting from the same datasets collected to generate Figure 2, which are commented in the previous section, we relabel the transitions with various sparse reward functions, one assigning reward of one to reach one specific state and zero to the others, repeated for every state of the environment. On those datasets, we train an agent offline and off-policy with Q-learning (Watkins & Dayan, 1992). At each training step, the agent samples a mini-batch of transitions $(s,a,r,s^{\prime})$ from the dataset and updates its Q-value as $Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ where $\alpha$ is the learning rate, $\gamma$ is the discount factor, and $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ represents the best estimated discounted future reward.
In Figure 3, we report the barplots of the corresponding offline RL success rate on each dataset relabeling, one for every possible goal state, ordering the results according to their respective success rate. Notably, a dataset collected with parallel agents allows to achieve the best offline RL performance across a large number of states, far greater than datasets collected with a single maximum entropy agent or a random policy. In the Appendix C, we provide further heatmaps visualizations showing the goal positions that offline RL algorithm is able to reach (Figure 9), and how each dataset is looking in terms of their empirical state distribution (Figure 8).
(a) Success Rate (ROOMDET)
(b) Success Rate   (LABDET)
(c) Success Rate (ROOMDET)
(d) Success Rate   (LABDET)
(e) Success Rate (ROOMDET)
(f) Success Rate     (LABDET)
@@ 7 Related Work
In the following, we summarize the most relevant works this paper sets its root in, spanning from entropy maximization to diverse skill discovery.
@@
State entropy maximization in MDPs has been introduced in Hazan et al. (2019), from which followed a variety of subsequent works focusing on the problem from various perspectives (Lee et al., 2019; Mutti & Restelli, 2020a; Mutti et al., 2021, 2022a, 2022b; Mutti, 2023; Zhang et al., 2021a; Guo et al., 2021; Liu & Abbeel, 2021b, a; Seo et al., 2021; Yarats et al., 2021; Nedergaard & Cook, 2022; Yang & Spaan, 2023; Tiapkin et al., 2023; Jain et al., 2023; Kim et al., 2023; Zisselman et al., 2023; Zamboni et al., 2024a, b).
@@
In the RL literature, first-order methods have been extensively employed to address non-concave policy optimization (Sutton et al., 1999; Peters & Schaal, 2008). In this work, we proposed a vanilla policy gradient estimator (Williams, 1992b) as a first step, yet several further refinements could be made, such as natural gradient (Kakade, 2001), trust-region schemes (Schulman et al., 2015), and importance sampling (Metelli et al., 2018).
@@
The problem of exploration in the presence of multiple agents is a vast and diverse topic. Among others, concurrent exploration  (Guo & Brunskill, 2015; Parisotto et al., 2019; Jung et al., 2020; Chen et al., 2022; Qin et al., 2022) has focused on scenarios where multiple agents operate within the same environment and learn simultaneously. Of particular interest are the works that investigate the role of information sharing between the agents (Alfredo & Arjun, 2017; Holen et al., 2023) or the ones that characterize the theoretical conditions for efficient coordinated exploration to happen and scale (Dimakopoulou & Van Roy, 2018; Dimakopoulou et al., 2018). Finally, exploration in multi-agent RL has gained attention but was almost limited to reward-shaping techniques based on many heuristics  Wang et al. (2019); Zhang et al. (2021b); Xu et al. (2024). In particular,Zhang et al. (2023) proposes a term maximizing the deviation from (jointly) explored regions.
@@
The problem of learning a set of diverse skills and/or policies is somehow linked to the problem of diversifying exploration via parallelization, and it has been addressed by a plethora of recent works (Hansen et al., 2019; Sharma et al., 2020; Campos et al., 2020; Liu & Abbeel, 2021a; He et al., 2022). Particularly relevant for this work are the results provided by Gregor et al. (2017); Eysenbach et al. (2019), in which the mutual information between visited states and skills is maximized, and the ones in Zahavy et al. (2022), that explicitly maximizes the diversity between policies.
@@ 8 Conclusions
This paper addressed the critical challenge of efficient exploration in reinforcement learning, particularly in scenarios with slow environment simulators or limited interaction budgets. We introduced a novel parallel learning framework that leverages the power of multiple agents to maximize state entropy while explicitly promoting diversity in their exploration strategies. Unlike traditional approaches that rely on multiple instances of identical agents, our method, PGPSE based on policy gradient techniques, employs a centralized learning mechanism that balances individual agent entropy with inter-agent diversity, effectively minimizing redundancy and maximizing the information gained from parallel exploration.
Our theoretical analysis, focusing on concentration properties, demonstrated the significant impact of parallelization on the rate of entropy stabilization and exploration diversity. Specifically, we showed that parallel agents, by focusing on different parts of the state space, can achieve faster convergence to high-entropy distributions compared to a single agent. These theoretical findings were strongly corroborated by our empirical results in various grid-world environments. PGPSE consistently outperformed single-agent baselines in terms of both state entropy and support size, demonstrating the effectiveness of our diversity-promoting approach.
Furthermore, we showed that the datasets collected by our parallel agents exhibit higher entropy and lead to improved performance in offline reinforcement learning settings, highlighting the practical value of our method for data-efficient learning. Although our results are promising, we acknowledge certain limitations. Our current framework is evaluated primarily on discrete grid-world environments. Future work will involve extending PGPSE to continuous and more complex environments, such as those found in robotics and control problems. Future work will extend the concept of state entropy maximization to the trajectory-based one, with centrally, the benefit of parallelizing the exploration strategy to achieve wider and diverse skills acquisition over complicated real-world problems. These extensions will be crucial for demonstrating the scalability of our approach to a broader use case.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Main Proofs and Additional Results
    The value of $p_{i}$ that minimizes $\frac{p_{i}(1-p_{i})}{\mathcal{H}^{2}(p_{i})}$ is the one with the highest entropy (see Figure4). Since the higher the entropy $\mathcal{H}(p_{i})$ the higher is also the variance $p_{i}(1-p_{i})$, we can restate the minimization problem as the ratio of two maximization problems.
@@
    The term at the numerator is the maximum variance, which can be lower bounded by the average variance.
@@
Frank-Wolfe for Parallel Exploration with a Infinite Trials
@@ Appendix B Experimental Details
In this paper, we present two distinct grid-world environments designed to illustrate the effectiveness of parallel exploration in enabling agents to overcome exploration bottlenecks and efficiently discover optimal trajectories in both structured and intricate scenarios.. In the first, two rooms are connected through a corridor with the goal positioned in the left room. In the second one, we considered a labyrinth-like space, with a more complex layout with many bifurcations but only one path leading to a goal. The two environments are shown in Figures 6 and 6. In the following, we briefly describe the two environments more thorough.
@@
Figure 6: Labirinth-like gridworld, with multiple path, goal state in the upper section
@@ B.1 ROOM
This is a $5\times 11$ grid world inspired by the toy example from (Towers et al., 2024). It consists of two rooms connected by a single horizontal corridor. The agent moves until reaching the goal or the episode terminates. The starting position is fixed at the center of the corridor, $(2,5)$. The agent navigates the grid using actions from $\{0,1,2,3\}$, where the index indicate 0: Move left, $1$: Move down, $2$: Move right, $3$: Move up. The observation space represents the player’s current position as an index computed by $\text{obs}=\text{row}\times\text{ncols}+\text{col}$ where rows and columns are zero-indexed. For instance, the goal position $(4,0)$ is mapped to $4\times 11+0=44.$. The total number of observations depends on the grid size. The environment is evaluated in both reward-free and sparse-reward settings, where in the latter case reward function $r:S\times A\to\mathbb{R}$ with $r=1$ upon reaching the goal and 0 otherwise. Throughout the paper, we consider two configurations:
  • ROOMDET: A deterministic version where executing an action $a$ under policy $\pi_{\theta}$ always results in the intended movement.
  • ROOMS: A stochastic variant where the agent plays the intended action with probability 1p1-p and deviate from it with $p$.
B.2 LAB

### v2

@@ 2 Agents1 Agent $K^{\prime}=2K$6 Agents1 Agent $K^{\prime}=6K$
(a) State Entropy (Room-det)
(b) State Entropy (Room-stoc)
(c) State Entropy (Maze-det)
(d) State Entropy (Maze-stoc)
(e) Support size (Room-det)
(f) Support size (Room-stoc)
(g) Support size (Maze-det)
(h) Support size (Maze-stoc)
@@
  • Policy Gradient Optimization: This part analyzes the results of the learning process based on Algorithm 5, evaluating the performance of the learned policy in terms of normalized entropy and support size.
  • Dataset Entropy Evaluation: This part takes a deeper look at the entropy of a dataset collected from the policies trained as in the previous section.
  • Offline RL: This part provides a comparative analysis of how the dataset collected with parallel maximum entropy agents benefits the performance of offline RL.
For all the experiments, we consider variations of discrete grid-world domains, which we depicted in Appendix B (see Figures 66). One family (referred to as Room) is composed of two rooms connected by a narrow corridor, with a total of 43 states and 4 actions (one to move in each direction). The other (referred to as Maze) is a maze of various connected paths, with same number of total states and actions as before. For each family, we consider a deterministic version (det) and a stochastic version (stoc) in which the agent’s chosen actions are flipped to random actions with some probability. Appendix C reports further experimental details, with additional visualizations and results.
@@
First of all, we test the ability of PGPSE (Algorithm 5) in addressing maximum state entropy exploration with parallel agents. We run multiple instances of the algorithm with an increasing number of agents $m\in\{2,4,6\}$. [^3] In the main text we omit the result with 4 agents, which are reported in Figure 7 of the Appendix. As a baseline, we consider the performance of the same algorithm training a single agent for the finite-trial objective (Eq. 2), with a number of trials $K^{\prime}$ matching the number of trajectories taken by the parallel agents. We compare the results in terms of (normalized) state entropy and the size of the entropy support (i.e., the number of unique visited states) in all the domains.
Figure 1 reports the resulting learning curves, obtained over five independent runs. For a fair comparison of algorithms with varying number of agents, the performance is reported as a function of the policy updates, which translate to $B\times m\times K\times T$ total interactions with the environment. Across the board, the performance of parallel agents (solid lines) consistently outperforms the single agent counterparts (dashed lines). Indeed, they achieve higher entropy values, thanks to their ability to explore more states as it is evident from the larger size of the support. This behavior can be attributed to the intrinsic diversity introduced by parallel agents: To optimize the team state distribution, each agent learns an individual behavior that is synergic with the others.
@@
Then, we tested the capabilities of the trained policies in collecting diverse experiences, by analyzing the properties of a single dataset. Figure 2 illustrates the normalized empirical state entropy of each dataset collected by the $m$ agents, compared against the entropy of the dataset collected by a single agent over $K=m$ trajectories (trained to maximize the entropy on the corresponding number of trajectories), and a random policy.
To construct these datasets, we sample one trajectory per agent from the parallel policies, evaluating them across five seeds. This procedure reveals that datasets generated through parallel exploration exhibit higher entropy than those collected by a single agent or a random policy. It is noteworthy that experience collected in parallel also exhibits minimal variance across runs, indicating that agents tend to generate trajectories in a quasi-deterministic manner, preserving consistency in the collected data.
2 Agents1 Agent $K^{\prime}\!\!\!=\!2K$6 Agents1 Agent $K^{\prime}\!\!\!=\!6K$Random
(a) Dataset Entropy (Room-det)
(b) Dataset Entropy (Room-stoc)
(c) Dataset Entropy (Maze-det)
(d) Dataset Entropy (Maze-stoc)
Figure 2: Entropy evaluation of the dataset generated by sampling from parallel policies, compared against both random and single-baselines across different environments. With the notation $K^{\prime}=mK$ we denoted the set in which the single agent had access to a multiple $m$ of trajectories available to the parallel agents. We report average and standard deviation over 5 independent runs.
(a) Success Rate, Parallel Dataset (Room-det)
(b) Success Rate, Parallel Dataset (Maze-det)
(c) Success Rate, Single Agent Dataset (Room-det)
(d) Success Rate, Single Agent Dataset (Maze-det)
(e) Success Rate, Random Dataset (Room-det)
(f) Success Rate, Random Dataset (Maze-det)
@@ 2 Agents1 Agent $K^{\prime}\!\!\!=\!2K$6 Agents1 Agent $K^{\prime}\!\!\!=\!6K$Random
Finally, we address the following question: Can datasets collected with parallel maximum entropy agents benefit offline RL algorithms? One practical application of the policies learn with PGPSE, in line with the motivation of this work, is to utilize the experience gained during the exploration phase to construct informative datasets for offline RL (Yarats et al., 2022).
Starting from the datasets collected as for Figure 2, we re-label the transitions with sparse reward functions, assigning rewards of +1+1 for a specific goal state and 0 otherwise. We then train an agent with offline Q-learning (Watkins & Dayan, 1992). At each training step, the agent samples a mini-batch of transitions $(s,a,r,s^{\prime})$ from the dataset and updates its Q-value as $Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ where $\alpha$ is the learning rate, $\gamma$ is the discount factor, and $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ represents the best estimated discounted future reward.
In Figure 3, we report the barplots of the corresponding offline RL success rate on each dataset relabeled, one for every possible goal state, ordering the results according to their respective success rate. Notably, a dataset collected with parallel agents allows to achieve the best offline RL performance across a large number of states, far greater than datasets collected with a single maximum entropy agent or a random policy. In Appendix C, we provide further heatmaps visualizations showing the goal positions that offline RL algorithm is able to reach (Figure 9), and how each dataset is looking in terms of their empirical state distribution (Figure 8).
Takeaways. Parallel State Entropy Maximization leads to better exploration, diverse data collection and improved performances in offline learning in PMDPs.
@@ 7 Related Work
In the following, we summarize the most relevant works this paper sets its root in, spanning from entropy maximization, multi-agent exploration to diverse skill discovery.
@@
State entropy maximization in MDPs has been introduced in Hazan et al. (2019), from which followed a variety of subsequent works focusing on the problem from various perspectives (Lee et al., 2019; Mutti & Restelli, 2020; Mutti et al., 2021, 2022a, 2022b; Mutti, 2023; Zhang et al., 2021a; Guo et al., 2021; Liu & Abbeel, 2021b, a; Seo et al., 2021; Yarats et al., 2021; Nedergaard & Cook, 2022; Yang & Spaan, 2023; Tiapkin et al., 2023; Jain et al., 2023; Kim et al., 2023; Zisselman et al., 2023; Zamboni et al., 2024a, b, 2025).
@@
In the RL literature, first-order methods have been extensively employed to address non-concave policy optimization (Sutton et al., 1999; Peters & Schaal, 2008). In this work, we proposed a vanilla policy gradient estimator (Williams, 1992) as a first step, yet several further refinements could be made, such as natural gradient (Kakade, 2001), trust-region schemes (Schulman et al., 2015), and importance sampling (Metelli et al., 2018).
@@
The problem of exploration in the presence of multiple agents is a vast and diverse topic. Among others, concurrent exploration  (Guo & Brunskill, 2015; Parisotto et al., 2019; Jung et al., 2020; Chen et al., 2022; Qin et al., 2022) has focused on scenarios where multiple agents operate within the same environment and learn simultaneously. Of particular interest are the works that investigate the role of information sharing between the agents (Alfredo & Arjun, 2017; Holen et al., 2023) or the ones that characterize the theoretical conditions for efficient coordinated exploration to happen and scale (Dimakopoulou & Van Roy, 2018; Dimakopoulou et al., 2018). Finally, exploration in Multi-Agent RL (MARL, Albrecht et al., 2024) has gained attention but was almost limited to reward-shaping techniques based on many heuristics  Wang et al. (2019); Zhang et al. (2021b); Xu et al. (2024). In particular, Zhang et al. (2023) proposes a term maximizing the deviation from (jointly) explored regions. Yet, we highlight a key distinction: In conventional MARL, multiple agents operate concurrently within a shared environment, requiring coordination due to the interdependence of their trajectories; In contrast, our setting involves multiple agents interacting with independent instances of the environment. While coordination remains beneficial for enabling specialization and aligning agents with distinct objectives, their trajectories are decoupled. This structural difference fundamentally separates our framework from typical MARL formulations like Li et al. (2021)Zhao et al. (2021) and Lupu et al. (2021).
@@
The problem of learning a set of diverse skills and/or policies is somehow linked to the problem of diversifying exploration via parallelization, and it has been addressed by a plethora of recent works (Hansen et al., 2019; Sharma et al., 2020; Campos et al., 2020; Liu & Abbeel, 2021a; He et al., 2022). The results provided by Gregor et al. (2017); Eysenbach et al. (2019) are particularly relevant for this work, in which the mutual information between visited states and skills is maximized, toghether with the ones in Zahavy et al. (2022), that explicitly maximizes the diversity between policies.
@@ 8 Conclusions
This paper targeted the critical challenge of efficient exploration in reinforcement learning, particularly in scenarios with slow environment simulators or limited interaction budgets. We introduced a novel parallel learning framework that leverages the power of multiple agents to maximize state entropy while explicitly promoting diversity in their exploration strategies. Unlike traditional approaches that rely on multiple instances of identical agents, our method, PGPSE based on policy gradient techniques, employs a centralized learning mechanism that balances individual agent entropy with inter-agent diversity, effectively minimizing redundancy and maximizing the information gained from parallel exploration.
Our theoretical analysis demonstrated the significant impact of parallelization on the rate of entropy stabilization and exploration diversity. Specifically, we showed that parallel agents, by focusing on different parts of the state space, can achieve faster convergence to high-entropy distributions compared to a single agent. These theoretical findings were strongly corroborated by our empirical results on various gridworld environments. PGPSE consistently outperformed single-agent baselines in terms of both state entropy and support size, demonstrating the effectiveness of our diversity-promoting approach.
Furthermore, we showed that the datasets collected by parallel agents exhibit higher entropy and lead to improved performance in offline settings, highlighting the practical value of our method for data-efficient learning. While our results are promising, we acknowledge certain limitations. Our current framework is evaluated primarily on discrete gridworld environments. Future work will involve extending PGPSE to continuous and more complex environments, such as those found in robotics and control problems.Future work will extend, the concept of state entropy maximization to the trajectory based one, with centrally, the benefit of parallelizing the exploration strategy to achieve wider and diverse skills acquisition over complicated real world problems. These extensions will be crucial for demonstrating the scalability of our approach to a broader use case.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Main Proofs and Additional Results
    The value of $p_{i}$ that minimizes $\frac{p_{i}(1-p_{i})}{\mathcal{H}^{2}(p_{i})}$ is the one with the highest entropy (see Figure4). The ration between the variance p(1p)p(1-p) and the squared entropy $\mathcal{H}^{2}(p)$ is symmetric about $p=\frac{1}{2}$ and it has negative derivative for $p<\frac{1}{2}$ and positive derivative for $p>\frac{1}{2}$. Since the higher the entropy $\mathcal{H}(p_{i})$, the higher is also the variance $p_{i}(1-p_{i})$, we can restate the minimization problem as the ratio of two maximization problems.
@@
    The term in the numerator is the maximum variance, which can be lower bounded by the average variance.
@@
Frank-Wolfe for Parallel Exploration with Infinite Trials
@@ Appendix B Experimental Details
In this paper, we present two distinct grid-world environments designed to illustrate the effectiveness of parallel exploration in enabling agents to overcome exploration bottlenecks and efficiently discover optimal trajectories in both structured and intricate scenarios.. In the first, two rooms are connected through a corridor with the goal positioned in the left room. In the second one, we considered a maze-like space, with a more complex layout with many bifurcations but only one path leading to a goal. The two environments are shown in Figures 6 and 6. In the following, we briefly describe the two environments more thorough.
@@
Figure 6: Maze gridworld, with multiple path, goal state in the upper section
@@ B.1 Room
This is a $5\times 11$ grid world inspired by the toy example from Towers et al. (2024). It consists of two rooms connected by a single horizontal corridor. The agent moves until reaching the goal or the episode terminates. The starting position is fixed at the center of the corridor, $(2,5)$. The agent navigates the grid using actions from $\{0,1,2,3\}$, where the index indicate 0: Move left, $1$: Move down, $2$: Move right, $3$: Move up. The observation space represents the player’s current position as an index computed by $\text{obs}=\text{row}\times\text{ncols}+\text{col}$ where rows and columns are zero-indexed. For instance, the goal position $(4,0)$ is mapped to $4\times 11+0=44.$. The total number of observations depends on the grid size. The environment is evaluated in both reward-free and sparse-reward settings, where in the latter case reward function $r:S\times A\to\mathbb{R}$ with $r=1$ upon reaching the goal and 0 otherwise. Throughout the paper, we consider two configurations:
  • Room-det: A deterministic version where executing an action $a$ under policy $\pi_{\theta}$ always results in the intended movement.
  • Room-stoc: A stochastic variant where the agent plays the intended action with probability 1p1-p and deviate from it with $p$.
B.2 Maze